OpenStudy (anonymous):

If f(x) = 9 − 2|x + 2|, find f(−7). −1 −15 1 19

OpenStudy (sheraz12345):

-1

OpenStudy (jojo4eva):

The answer is -1

OpenStudy (jojo4eva):

sustitute the -7 in for x -7+2=-5 which turns into 5 because of the absolute value then you multiply first which 2*5=10 the 9-10=-1

OpenStudy (anonymous):

thanks guys

OpenStudy (jojo4eva):

no probs